Calculate the mass in grams of NaOH that must be added to make a 15% solution. Density of water= 1g/ml.

1
Expert's answer
2021-12-13T17:03:03-0500

Mass percentage "(\frac{W}{W})" is the mass of solute ( in grams) in "100g" solution expressed in percentage


Mass by volume percentage "(\frac{W}{V})" is the mass of solute present in "100ml" of aqueous solution.


The "15\%" solution of NaOH is not specified.

Assuming it is "15\%" in reference to mass percentage "(\frac{W}{W});"


"15g" of NaOH will be dissolved in "(100-15)=85g" of water


Volume of "85g" of water "=\frac{85}{1}=85ml"


Therefore "15g" of NaOH will be dissolved in 85ml of water.


From mass by volume percentage concept;

"15g" of NaOH will be dissolved to make 100ml of solution